# Joint Estimation and Prediction of City-wide Delivery Demand: A Large Language Model Empowered Graph-based Learning Approach

### v1

@@ 3.2.2 Parameterization by Space-then-Time Message Passing Network
To simplify the process, we assume the relational structure between nodes keeps static, and the above equation can be decoupled and instantiated as a time-then-space architecture (Gao & Ribeiro, 2022):
@@ 4.1.1 Location-based Geospatial Query
The sensory data such as demand series is highly associated with some location-dependent variables. However, collecting sufficient geospatial information, including points of interest, road network geometry, and buildings, can be a laborious and tedious process. This is especially difficult when different cities have different sources and availability of location data. Fortunately, recent advances in natural language processing provide an alternative paradigm for mining geospatial knowledge (Gurnee & Tegmark, 2023; Manvi et al., 2024). As LLMs have shown an impressive ability to reason about geospatial data (Manvi et al., 2023), it is expected to extract general location-based representations based on the intrinsic knowledge of LLMs.
To achieve this, we first construct a location-based query and organize a text prompt as input to LLMs. First, given the query coordinate of the target location, a public map service or database such as OpenStreetMap (OSM) (Haklay & Weber, 2008) is accessed and some general geolocation information is obtained, such as POIs and surrounding streets and buildings. Then a text message containing the basic location information extracted from the OSM is generated as a prompt and fed into LLMs, as demonstrated in Fig. 2. LLMs are requested to identify and understand the geolocation information, relying on their ability to reason in context.
@@ 4.1.2 Prompt-based Geospatial Reasoning and Encoding Generation
Different from existing frameworks that adopt LLMs as the backbone forecasting backbone (Jin et al., 2023a), or directly generate geospatial message from the output (Manvi et al., 2023), we propose to adopt the pre-activated embedding from the output layer of a pre-trained LLM as the latent covariate, which is assumed to be geospatially encoded by the network processing.
To extract the general geospatial encoding from LLMs that is applicable for downstream tasks, we use the linear probing technique (Alain & Bengio, 2016; Belinkov, 2022). Specifically, given the network activation of hidden states (before the last softmax function) of LLMs, we can use a linear ridge regressor as the probe to fit the node embedding of each ROI:
@@ 4.1.3 Functional Graph Construction with LLM Encoding
As discussed in section 3.2.1, the correlation and interaction between regions are described by a graph. Due to the absence of a precise metric to construct such a graph, physical proximity is generally adopted (Lei et al., 2022; Nie et al., 2023b). However, proximity in the Euclidean is inadequate to reflect the relationships of demand patterns in feature space. Fortunately, the inherent geospatial knowledge in the LLM-based encoding provides an abstract of the geolocation property of each region in the latent space. We can measure the similarity of these encodings and treat it as a more accurate surrogate of the functional similarity. To achieve this, the functional graph is learned end-to-end simultaneously with the main architecture:
@@ 4.3 End-to-End Transferable Predictor by Inductive Learning
Using the model defined above, a straightforward solution consists of adopting an estimation model to first estimate the demand for unobserved regions and then performing downstream forecasting processing. However, we argue that the future dynamics of demand evolution in existing regions (or depots) can be substantially influenced by the newly developed regions, showing a synchronous effect for their individual demand behaviors. Therefore, we propose bypassing the estimation step and considering the forecasting backbone that can directly handle irregularly structured graph data. We aim to jointly estimate demand of newly developed regions and predict the future demand values of already existing depots in an end-to-end way, benefiting from the optimization of the shared target.
This can be modeled as an inductive learning task (Wu et al., 2021b; c; Nie et al., 2023c) associated with a kriging problem (Appleby et al., 2020; Lei et al., 2022; Nie et al., 2023b). We first introduce an indicating variable to denote whether the region has records or is masked for training purposes. The inductive training scheme constructs a random masking and reconstruction task using the observed data. We then consider the following conditional probability during training:
@@
Based on above discussions, the model needs to address the forecasting problem with irregular graph structures, i.e., the testing graph and node locations can differ from the ones in training. We approach this by applying the masking-reconstruction scheme proposed in (Wu et al., 2021b). As illustrated in Fig. 4, this process includes random masking and reconstruction during the training stage. Consider that we can split the training data into context-target window pairs along the temporal dimension, the following scheme is established. , we have:
@@ 5.1 Dataset Description
The entire area of each city is divided into several geospatial regions for logistics management. The datasets provide event-specific information, such as the order ID, pickup time, location, GPS points, courier information, delivery event, etc. Each order is attached to the corresponding region of interests (ROIs) according to its pickup GPS record. For package data, we set a time interval of 1 hour and aggregated the orders within this specified window. For food data, due to the sparsity of this data record, we select a 12-hour interval. In both datasets, we calculate the total number of pickups, i.e., the pickup demand. In the following sections, we adopt both package and food delivery data for model comparisons. For subsequent model analysis and discussion, we mainly adopt the package delivery data for experiments. Total number of regions for each city is summarized in Tab. 2.
@@ 5.2 Experiment Settings
Baseline models. Apart from the historical average (HA) method, we select several representative models from highly related literature in machine learning studies, including: diffusion convolutional recurrent neural network (DCRNN) (Li et al., 2017), spatial-temporal convolutional neural network (Yu et al., 2017), graph wavenet (GWNET) (Wu et al., 2019a), graph learning neural network (MTGNN) (Wu et al., 2020), inductive graph neural network (IGNNK) (Wu et al., 2021a), spatial aggregation and temporal convolutional network (SATCN) (Wu et al., 2021c), message-passing gated recurrent unit (MPGRU) (Gao & Ribeiro, 2022), and graph recurrent imputation network (GRIN) (Cini et al., 2021). Particularly, DCRNN, STGCN, GWNET, MTGNN are originally designed for forecasting tasks; While IGNNK, SATCN, MPGRU, and GRIN are designed for kriging and imputation problems. We evaluate both of the two categories for a fair comparison.
Tasks setups. The models are trained to predict the future demand of all locations in the next 24 steps, given a 24-step historical window of existing regions. For Shanghai, Hangzhou, Chongqing and Yantai, we randomly choose 10 regions as newly developed regions for testing. For Jilin, Los Angeles, New York and San Francisco, the number of new regions is set to 5. It is noted that the unknown testing regions are unavailable for all models during training and only accessible for testing purposes. Following Wu et al. (2023), we use the 6:2:2 ratio for training, evaluation, and testing sets based on the chronological order of the timestamps. To assess the performance of the above models, we report the metrics of Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE).
@@ 5.3.2 Task 2: Model Transfers to New Cities with Full Observations
5.3.3 Task 3: Model Transfers to New Cities with New Depots
@@
When there exist regions without any historical record in both source and target domains, the challenge of model transfer is further intensified. Tab. 6 examines the performance of a mode trained in Shanghai and transferred to new cities with different numbers of new regions. As expected, the prediction error is larger than the results in Tab. 5 due to the existence of unobserved regions in both the training set and testing set. In this scenario, our model displays more significant superiority over baselines. This can be ascribed to the generality of LLM-based location encoding that has comprehensive and fundamental geospatial knowledge shared across different cities.
@@ 6.1 Understanding the Encoding from LLMs
An intriguing question pertains to the manner in which LLMs comprehend the geospatial relationship. To initiate an investigation into this matter, we present in Fig. 7 a visualization of the encoding produced by LLMs, employing the t-SNE dimensional reduction technique. Interestingly, the encoding can differentiate different cities, and regions belonging to the same city are aggregated closely. Cities in different countries are relatively far apart, which is in line with human understanding.
@@ 6.2.1 Ablation Study
w adjacency graph: we replace the LLM-based graph with a proximity-based graph;
@@ 6.2.2 Sensitivity Analysis
As an initial evaluation, Tab. 7 report the performances of separate estimation and forecasting tasks. These models are trained on respective tasks with separate objectives. It can be observed that Impel has a beneficial effect on the performance of both tasks, with a particularly notable improvement in the more difficult estimation task. Tab. 8 further sheds light on the necessity of the joint estimation and prediction framework. The separate treatment can achieve a lower training error by learning to fit the full distribution. However, they over-fit the training data and generate poor generalization performance in testing set where new regions are included. Instead, our end-to-end setting balances the capacity and robustness, achieving the best accuracy in this hybrid task.
@@ Acknowledgments
This research was sponsored by the National Natural Science Foundation of China (52125208), the National Natural Science Foundation of China’s Young Scientists Fund (52302413), the Science and Technology Commission of Shanghai Municipality (No. 22dz1203200), and the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU/25209221 and PolyU/15206322).

### v3

@@ 3.2.2 Parameterization by Space-then-Time Message Passing Network
To simplify the process, we assume the relational structure between nodes keeps static, and the above equation can be decoupled and instantiated as a time-then-space architecture (Gao and Ribeiro, 2022):
@@ 4.1.1 Location-based Geospatial Query
The sensory data such as demand series is highly associated with some location-dependent variables. However, collecting sufficient geospatial information, including points of interest, road network geometry, and buildings, can be a laborious and tedious process. This is especially difficult when different cities have different sources and availability of location data. Fortunately, recent advances in natural language processing provide an alternative paradigm for mining geospatial knowledge (Gurnee and Tegmark, 2023; Manvi et al., 2024). As LLMs have shown an impressive ability to reason about geospatial data (Manvi et al., 2023), it is expected to extract general location-based representations based on the intrinsic knowledge of LLMs.
To achieve this, we first construct a location-based query and organize a text prompt as input to LLMs. First, given the query coordinate of the target location, a public map service or database such as OpenStreetMap (OSM) (Haklay and Weber, 2008) is accessed and some general geolocation information is obtained, such as POIs and surrounding streets and buildings. Then a text message containing the basic location information extracted from the OSM is generated as a prompt and fed into LLMs, as shown in Fig. 2. LLMs are requested to identify and understand the geolocation information, relying on their ability to reason in context.
@@ 4.1.2 Prompt-based Geospatial Reasoning and Encoding Generation
Different from existing frameworks that adopt LLMs as the backbone forecasting backbone (Jin et al., 2023a), or directly generate geospatial message from the output (Manvi et al., 2023), we propose to adopt the pre-activated embedding from the output layer of a pre-trained LLM as the latent covariate, which is assumed to be geospatially encoded by the network processing. To extract the general geospatial encoding from LLMs that is applicable for downstream tasks, we use the linear probing technique (Alain and Bengio, 2016; Belinkov, 2022). Specifically, given the network activation of hidden states (before the last softmax function) of LLMs, we can use a linear ridge regressor as the probe to fit the node embedding of each ROI:
@@ 4.1.3 Functional Graph Construction with LLM Encoding
As discussed in section 3.2.1, the correlation and interaction between regions are described by a graph. Due to the absence of a precise metric to construct such a graph, physical proximity is generally adopted (Lei et al., 2022; Nie et al., 2023a). However, proximity in the Euclidean is inadequate to reflect the relationships of demand patterns in feature space. Fortunately, the inherent geospatial knowledge in the LLM-based encoding provides an abstract of the geolocation property of each region in the latent space. We can measure the similarity of these encodings and treat it as a more accurate surrogate of the functional similarity. To achieve this, the functional graph is learned end-to-end simultaneously with the main architecture:
@@ 4.3 End-to-End Transferable Predictor by Inductive Learning
Using the model defined above, a straightforward solution consists of adopting an estimation model to first estimate the demand for unobserved regions and then performing downstream forecasting processing. However, we argue that the future dynamics of demand evolution in existing regions (or depots) can be substantially influenced by the newly developed regions, showing a synchronous effect for their individual demand behaviors. Therefore, we propose bypassing the estimation step and considering the forecasting backbone that can directly handle irregularly structured graph data. We aim to jointly estimate demand of newly developed regions and predict the future demand values of already existing regions in an end-to-end way, benefiting from the optimization of the shared target.
This can be modeled as an inductive learning task (Wu et al., 2021a, c; Nie et al., 2023b) associated with a kriging problem (Appleby et al., 2020; Lei et al., 2022; Nie et al., 2023a). We first introduce an indicating variable to denote whether the region has records or is masked for training purposes. The inductive training scheme constructs a random masking and reconstruction task using the observed data. We then consider the following conditional probability during training:
@@
Based on above discussions, the model needs to address the forecasting problem with irregular graph structures, i.e., the testing graph and node locations can differ from the ones in training. We approach this by applying the masking-reconstruction scheme proposed in (Wu et al., 2021a). As illustrated in Fig. 4, this process includes random masking and reconstruction during the training stage. Consider that we can split the training data into context-target window pairs along the temporal dimension, the following scheme is established. , we have:
@@ 5.1 Dataset Description
The entire area of each city is divided into several geospatial regions for logistics management. The datasets provide event-specific information, such as the order ID, pickup time, location, GPS points, courier information, delivery event, etc. Each order is attached to the corresponding region of interests (ROIs) according to its pickup GPS record. For package data, we set a time interval of 1 hour and aggregated the orders within this specified window. For food data, due to the sparsity of this data record, we select a 12-hour interval. In both datasets, we calculate the total number of pickups, i.e., the pickup demand. In following sections, we adopt both package and food delivery data for model comparisons. Individual orders are anonymized and aggregated into regions to eliminate privacy breaches. For subsequent model analysis and discussion, we mainly adopt the package delivery data for experiments. Total number of regions for each city is summarized in Tab. 2.
@@ 5.2 Experiment Settings
Baseline models. Apart from the historical average (HA) method, we select several representative models from highly related literature in machine learning studies, including: diffusion convolutional recurrent neural network (DCRNN) (Li et al., 2017), spatial-temporal convolutional neural network (Yu et al., 2017), graph wavenet (GWNET) (Wu et al., 2019b), graph learning neural network (MTGNN) (Wu et al., 2020), inductive graph neural network (IGNNK) (Wu et al., 2021b), spatial aggregation and temporal convolutional network (SATCN) (Wu et al., 2021c), message-passing gated recurrent unit (MPGRU) (Gao and Ribeiro, 2022), and graph recurrent imputation network (GRIN) (Cini et al., 2021). Particularly, DCRNN, STGCN, GWNET, MTGNN are originally designed for forecasting tasks; While IGNNK, SATCN, MPGRU, and GRIN are designed for kriging and imputation problems. We evaluate both of the two categories for a fair comparison.
Tasks setups. Models are trained to predict the future demand of all locations in the next 24 steps, given a 24-step historical window of existing regions. For Shanghai, Hangzhou, Chongqing and Yantai, we randomly choose 10 regions as newly developed regions for test. For Jilin, Los Angeles, New York and San Francisco, the number of new regions is set to 5. Note that the unknown testing regions are unavailable for all models during training and only accessible for testing purposes. Following Wu et al. (2023), we use the 6:2:2 ratio for training, evaluation and test sets based on the chronological order of the timestamps. To assess the performance of the above models, we report the metrics of Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE).
@@ 5.3.2 Task 2: Model Transfers to New Cities with Full Observations
5.3.3 Task 3: Model Transfers to New Cities with New Regions
@@
When there exist regions without historical data in both source and target domains, the challenge of model transfer is further intensified. Tab. 6 examines the performance of a mode trained in Shanghai and transferred to new cities with different numbers of new regions. As expected, the prediction error is larger than the results in Tab. 5 due to the existence of unobserved regions in both the training set and the test set. In this scenario, our model shows more significant superiority over baselines. This can be ascribed to the generality of LLM-based location encoding that has comprehensive and fundamental geospatial knowledge shared across cities.
@@ 6 Discussions
6.1 Benchmarking the LLM-based Geolocation Encoding
To justify the significance of LLM-based geolocation encoding in delivery modeling, we perform experiments to compare it with other geo-encoding methods. This section contains the following evaluations:
- 1.
2.
Encoding POIs with large embedding models, such as the General Text Embeddings (GTE) (Li et al., 2023);
3.
Using one-hot encoding to process POI features;
4.
Only encoding of target address with LLMs.
| Models | Shanghai | Hangzhou | Chongqing | Jilin | Yantai | |||||
| MAE | RMSE | MAE | RMSE | MAE | RMSE | MAE | RMSE | MAE | RMSE | |
| Space2Vec | 4.28 | 7.93 | 5.93 | 10.75 | 6.47 | 10.28 | 1.52 | 2.68 | 3.41 | 6.61 |
| Sphere2Vec | 4.33 | 8.23 | 10.26 | 20.17 | 3.13 | 5.54 | 1.47 | 2.67 | 3.55 | 6.73 |
| LLM+Address | 5.95 | 14.17 | 4.99 | 10.48 | 2.96 | 5.38 | 1.98 | 4.08 | 2.94 | 5.22 |
| POI Only | 6.98 | 16.00 | 11.44 | 22.07 | 7.67 | 12.03 | 2.69 | 5.48 | 4.06 | 7.79 |
| GTE+POI | 4.08 | 8.71 | 5.32 | 11.88 | 3.04 | 5.82 | 1.90 | 4.05 | 2.89 | 4.86 |
Results are shown in Table 7. As can be seen, our model consistently achieves the best performances. Pretrained geographic methods such as Space2Vec cannot provide effective contextual information for delivery modeling. Additionally, encoding only the coordinates (address) or using other methods to process POIs cannot achieve desirable performances.
6.2 The generality of LLM-based Geolocation Encoding in Enhancing Demand Prediction
Recall that our model has a simple-yet-effective architecture: it bases on stacked spatiotemporal message passing layers that are easy to compute. This efficiency is achieved by the effectiveness of LLM-based encoding. To evaluate whether the enhancement of LLM-based encoding is generic, we equip other baselines with the same encoding and report their performances. The results presented in Tab. 8 indicate that all baseline architectures can benefit from the embedding, with improvements exceeding 10%. This result means that better prediction accuracy can be achieved by adopting more advanced backbone architectures.
| Models | Shanghai | Hangzhou | IMP | ||
| MAE | RMSE | MAE | RMSE | (%) | |
| DCRNN | 5.65 | 11.86 | 7.33 | 14.59 | 14.83% |
| +LLM Enc. | 4.86 | 9.98 | 6.34 | 12.55 | |
| STGCN | 5.07 | 11.62 | 6.38 | 14.26 | 18.64% |
| +LLM Enc. | 4.47 | 9.34 | 5.36 | 10.39 | |
| GWNET | 5.22 | 11.67 | 7.99 | 15.90 | 21.04% |
| +LLM Enc. | 4.58 | 9.25 | 6.03 | 11.69 | |
| MTGNN | 5.09 | 11.56 | 6.23 | 13.89 | 21.92% |
| +LLM Enc. | 4.05 | 8.15 | 5.45 | 10.38 | |
| IGNNK | 5.22 | 11.50 | 7.25 | 15.06 | 25.52% |
| +LLM Enc. | 3.95 | 7.90 | 5.88 | 11.21 | |
| SATCN | 4.75 | 9.38 | 7.64 | 14.77 | 13.68% |
| +LLM Enc. | 4.58 | 8.85 | 6.12 | 11.00 | |
@@ 6.3 Understanding the Encoding from LLMs
An intriguing research question pertains to the manner in which LLMs comprehend the geospatial relationship. To initiate an investigation into this matter, we present in Fig. 7 a visualization of the encoding produced by LLMs, employing the t-SNE dimensional reduction technique. Interestingly, the encoding can differentiate different cities, and regions belonging to the same city are aggregated closely. Cities in different countries are relatively far apart, which is in line with human understanding.
@@ 6.4.1 Ablation Study
w adjacency graph: we replace the LLM-based graph with a proximity (distance)-based graph;
@@ 6.4.2 Sensitivity Analysis
As an initial evaluation, Tab. 9 report the performance of separate estimation and forecasting tasks. These models are trained on respective tasks with separate objectives. It can be seen that Impel has a beneficial effect on the performance of both tasks, with a particularly notable improvement in the more difficult estimation task. Tab. 10 further sheds light on the necessity of the joint estimation and prediction framework. Separate treatment can achieve a lower training error by learning to fit the full distribution. However, they over-fit the training data and generate poor generalization performance in testing set where new regions are included. Instead, our end-to-end setting balances the capacity and robustness, achieving the highest accuracy in this hybrid task.
@@ 7 Conclusion and Future Work
Although promising results have been demonstrated, there are several shortcomings need further attention. First, it is important to ensure the interpretability of the model within the logistics and delivery sector to maintain the trust of stakeholders. Our model is currently based on black-box architectures in which explainable AI techniques can be used to interpret it. Second, the estimation and transfer of demand in new regions are based on the observation of existing regions, which can lead to error accumulation.
@@ Acknowledgments
This research was sponsored by the National Natural Science Foundation of China (52125208), the National Natural Science Foundation of China’s Fundamental Research Program for Young Students (524B2164), and grants from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU/15206322 and PolyU/15227424).
8 Appendix
8.1 More explanations on the role of LLM-based encoding
We provide more explanations on the role of LLM-derived embedding in the context of demand prediction and estimation. Actually, the demand prediction problem (without missing observations) can generally be considered as a time series forecasting task in which deep neural networks are applied to predict future value based on historical data. In state-of-the-art time series forecasting models, three types of additional features can be collected and used as covariates (Cini et al., 2023b): (1) time-related features, such as time-of-day and day-of-week encodings; (2) region-specific variates, such as the region (sensor) identity information and property; (3) cross-region variates, such as region-wide relationships described by an adjacency graph. These features complement the historical-future data pairs, and this setting is commonly acknowledged in previous work (Li et al., 2017; Wu et al., 2019b; Yu et al., 2017; Cini et al., 2021; Wu et al., 2020; Nie et al., 2024a).
For (1), we have performed supplementary experiments by adding timestamp features to the model in subsection 8.2.2, and the results show that the additional temporal features do not benefit the performances significantly; For (2), we use the LLM-based location encoding for each region as specialized features; For (3), we construct a functional graph based on the LLM-based encoding to quantify the relationship between regions. Therefore, the LLM-derived embedding can characterize rich information to improve prediction performance.
8.2 Supplementary experiments
8.2.1 Experiments on the observation patterns
It should be noted that the experiments described in section 5 are based on the random selection and masking of regions for evaluation purposes. This may not represent the real-world scenarios of service area expansion. To mitigate this experimental bias, we perform additional experiments that remove a whole cluster of regions that are close to each other, which leads to a cluster missing pattern. Note that the random seed is fixed for all baselines for the same city for a fair comparison. Tab. 11 provides the results on package delivery datasets. As expected, most of the baselines show worse performance due to the difficulty in model generalization. Interestingly, in some cases our model performs better than the random missing scenario and has a greater margin of improvement. This is due to the effectiveness of the LLM-based encoding for providing sufficient geolocation information about the target location.
| Models | Shanghai | Hangzhou | Chongqing | Yantai | ||||
| MAE | RMSE | MAE | RMSE | MAE | RMSE | MAE | RMSE | |
| DCRNN (Li et al., 2017) | 5.77 | 12.30 | 6.32 | 11.91 | 4.42 | 8.36 | 2.65 | 5.05 |
| STGCN (Yu et al., 2017) | 5.60 | 12.86 | 6.33 | 10.98 | 4.52 | 10.71 | 2.61 | 5.68 |
| GWNET (Wu et al., 2019b) | 13.06 | 25.21 | 6.32 | 11.43 | 4.52 | 10.24 | 2.72 | 5.73 |
| MTGNN (Wu et al., 2020) | 5.56 | 12.50 | 6.09 | 11.37 | 4.57 | 10.90 | 2.77 | 5.72 |
| IGNNK (Wu et al., 2021b) | 5.86 | 13.21 | 5.82 | 10.44 | 4.72 | 10.89 | 3.11 | 6.76 |
| SATCN (Wu et al., 2021c) | 5.06 | 9.31 | 6.57 | 11.15 | 3.64 | 6.62 | 3.37 | 5.78 |
| MPGRU (Gao and Ribeiro, 2022) | 6.64 | 14.55 | 7.24 | 13.11 | 5.35 | 11.89 | 3.51 | 7.46 |
| GRIN (Cini et al., 2021) | 5.55 | 12.76 | 5.46 | 10.36 | 4.67 | 11.14 | 2.85 | 6.19 |
| Impel (Ours) | 3.45 | 7.00 | 4.41 | 9.56 | 2.50 | 5.21 | 2.30 | 4.73 |
| Improvement () | 31.82 | 24.81 | 19.23 | 7.72 | 31.32 | 21.30 | 11.88 | 6.34 |
8.2.2 Experiments on the Temporal Modeling
A common practice in deep traffic prediction model is to process the sequence with temporal modules such as recurrent neural networks and timestamp features. Therefore, we examine the effect of temporal techniques in our architecture by equipping it with temporal modules and timestamp features. For timestamp features, we use the time-of-day and day-of-week positional encodings used in previous work (Cini et al., 2021) and concatenate them with the input time series. For additional temporal modules including RNN and TCN, we replace the original MLP in Eq. 25 with them. We report both the accuracy and training efficiency in Table 12. Performances of several baselines are also shown as a reference. It is observed that the gain of temporal modeling is marginal. However, introducing complicated temporal techniques has increased the computational burden. To this end, we choose our architecture for simplicity and better transferability.
| Models | Shanghai | Hangzhou | ||||||||
| MAE | RMSE | Memory | Speed | Params. | MAE | RMSE | Memory | Speed | Params. | |
| Impel (Ours) | 3.76 | 7.93 | 1.5 GB | 1.1 s/batch | 218.6 k | 4.52 | 9.90 | 1.6 GB | 1.2 s/batch | 218.6 k |
| DCRNN | 5.65 | 11.86 | 2.3 GB | 929.7 s/batch | 2215.8 k | 7.33 | 14.59 | 2.5 GB | 1311.8 s/batch | 2215.8 k |
| GWNET | 5.22 | 11.67 | 3.0 GB | 165.8 s/batch | 2434.7 k | 7.99 | 15.90 | 3.2 GB | 175.0 s/batch | 2435.8 k |
| MTGNN | 5.09 | 11.56 | 1.7 GB | 52.8 s/batch | 2275.3 k | 6.23 | 13.89 | 1.8 GB | 36.9 s/batch | 2338.1 k |
| w/ RNN | 4.38 | 8.87 | 1.8 GB | 8.8 s/batch | 483.1 k | 5.77 | 10.90 | 1.8 GB | 9.2 s/batch | 483.1 k |
| w/ TCN | 4.58 | 8.93 | 1.7 GB | 5.0 s/batch | 376.0 k | 4.89 | 10.75 | 1.8 GB | 5.3 s/batch | 376.0 k |
| w/ Time Feat. | 3.76 | 7.93 | 1.5 GB | 1.1 s/batch | 219.8 k | 4.50 | 9.90 | 1.6 GB | 1.2 s/batch | 219.8 k |
| Input 36 Steps | 3.73 | 7.74 | 1.5 GB | 1.1 s/batch | 219.4 k | 4.52 | 9.87 | 1.6 GB | 1.2 s/batch | 219.4 k |
| Input 48 Steps | 3.74 | 7.89 | 1.5 GB | 1.2 s/batch | 220.2 k | 4.55 | 9.95 | 1.6 GB | 1.4 s/batch | 220.2 k |
| Input 72 Steps | 3.77 | 7.97 | 1.6 GB | 1.4 s/batch | 221.7 k | 4.53 | 9.88 | 1.7 GB | 1.4 s/batch | 221.7 k |
8.2.3 Online Prediction Performances and Updating Strategy
It is important to consider a real-world scenario in which data is accumulated over time. As a result, the model will likely require regular updates and maintenance in order to maintain its accuracy. We are interested in how the model will be managed and updated with new data. Accordingly, we utilize a single pretrained model derived from a limited number of observed regions and deploy it to online prediction tasks involving an increasing number of novel regions. Fig. 11 shows the online prediction results in two cities. It is noted that this figure is different from the result in Fig. 9 as only a single model is applied. Our model shows better robustness than the baseline model IGNNK in the online application stage.
Since the model is trained in an inductive manner, it can deal with increasing data scales and region numbers. The requisite accuracy can be achieved through the implementation of model fine-tuning or retraining. Additionally, the LLM-based encoding can be precalculated for all possible regions offline. This makes the online application efficient.
Regarding the online test efficiency, we evaluate the computational efficiency at both inference stage (with generated LLM-based encoding) and the LLM encoding stage. Table 13 reports the total running time and memory consumption at two different stages. It is observed that the encoding stage requires a high computational burden due to the high complexity of LLMs. Fortunately, this encoding step (a single forward pass) only needs to be performed once throughout the testing or training phase and can also be performed offline after the location of the test points has been determined. After equipping the generated encoding into the prediction model, it can complete the testing task at low cost, preserving the capability of the online application.
| Models | Testing stage | LLM encoding stage | |||
| Running Time | GPU Memory | GPU Memory | Encoding Speed | Batch Size | |
| Impel | 0.17 s | 1.5 GB | 15.3 GB | 2.48 s | 1 |
| 15.5 GB | 2.04 s | 4 | |||
| 15.8 GB | 2.11 s | 8 | |||